Grade 11Physical Chemistry

calculate the empirical formula of gold chloride which contains 35.1% of chlorine.[atomic mass of Au=197]

kavya sharma
9 Years agoGrade 11

2 Answers

varun
ApprovedApproved Tutor Answer9 Years ago
HELLO KAVYA,
we are given that a compound conatins chlorine = 35.1%
and the rest will be gold
therefore,
100gm of a compound will contain = 100-35.1
                                                      =64.9%
molar mass of chlorine = 35.5g
so,number of moles in 35.1gm of chlorine = given mass/molar mass
                                                                   =35.1/35.5
                                                                   = 0.988 moles
Molar mass of gold = 197gm
number of moles in 64.9 gm of gold = given mass/molar mass                                                           = 64.9/197 
                                                            = 0.329 moles
NOW,we will calculate the simplest ratio of the elements by dividing the number of moles of each element by smallest value 0.329
Simplest ratio = Au(0.329/0.3289) Cl(0.988/0.329)
HENCE THE EMPIRICAL FORMULA OF THE COMPOUND IS AuCl3
                                               
Chitra
6 Years ago
100 gm of compound